OpenStudy (anonymous):

this is kind of a dumb question but how do you expand this? 3(a+h)²

OpenStudy (anonymous):

ooo easy!!!! come on girl u know u got this its.... 3(a+h)(a+h) none is a dumb question :)

OpenStudy (anonymous):

I wasnt sure if it became (3a+3h)² lol I should trust PEMDAS more.. thank you so much @lilly21

OpenStudy (anonymous):

Does this then become 3a²+3ah+3ah+3h²

OpenStudy (unklerhaukus):

\[3(a+h)^2=3(a+h)(a+h)=3(a^2+2ah+h^2)=3a^2+6ah+3h^2 \]

OpenStudy (anonymous):

Thank you @UnkleRhaukus you left me with no more questions left to ask! :)